Ann's age is 7/10 of Bill's age. In 4 years Ann will be 8/11 of Bill's age. How old is Bill now?
Answer by ankor@dixie-net.com(22740)

A = Ann's age now; B = Bill's age now
:
Just write what it says:
"Ann's age is 7/10 of Bill's age."
A = .7B
:
"In 4 years Ann will be 8/11 of Bill's age."
(A+4) = (8/11)(B+4)
:
Substitute .7B for A and you have:
.7B+4 = (8/11)(B+4)
:
Multiply equation by 11:
11(.7B+4) = (11)(8/11)(B+4)
:
and you have:
7.7B + 44 = 8(B+4)
:
7.7B + 44 = 8B + 32
;
44 - 32 = 8B - 7.7B
:
.3B = 12
:
How old is Bill now?
B = 12/.3
:
B = 40 yrs old
:
:
Check: that makes A = .7*40 = 28 yrs old
Using the statement, "In 4 years Ann will be 8/11 of Bill's age."
28 + 4 = 8/11(40 + 4)
32 = 8/11(44); proves our solution
:
Did this help?